# Data-Based System Representation and Synchronization for Multiagent Systems

### v1

@@ Abstract
This paper presents novel solutions of the data-based synchronization problem for continuous-time multiagent systems. We consider the cases of homogeneous and heterogeneous systems. First, a data-based representation of the synchronization error dynamics is obtained for homogeneous systems, using input-state data collected from the agents. Then, we show how to extend existing data-based stabilization results to the multiagent case to stabilize the obtained synchronization errors. The proposed method relies on the solution of a set of linear matrix inequalities that are shown to be feasible. Then, we solve the synchronization problem for heterogeneous systems by means of dynamic controllers. Different from existing results, we do not require model knowledge for the followers and the leader. The theoretical results are finally validated using numerical simulations.
@@ I INTRODUCTION
In recent years, there has been a great interest in developing stabilizing controllers without the need of any knowledge of a system’s model. One of the main drivers of the current data-based control research is a result now commonly known as Willems’ fundamental lemma [5], which provides a method to represent every input-output trajectory of a discrete-time linear system using only measured data. Continuous-time versions of this result have also been proposed [6, 7]. Willems’ lemma has been recently leveraged for the data-based design of stabilizing, robust, optimal and/or predictive controllers, as summarized in the survey [8] for discrete time systems. For examples of data-based control of continuous-time systems, see [9, 10, 11].
Some results have been obtained for the data-driven control of multiagent systems. In [12] the synchronization problem is solved using only partial knowledge of the agents’ dynamics. In [13], consensus protocols are designed for systems with identical models using time-varying feedback matrices. Several iterative learning algorithms have been proposed to achieve synchronization as in, e.g., [14, 15]. In [16], a different setting is considered as a multiagent predictive control scheme is used to minimize a global cost function.
In this paper, we study the data-based synchronization problem for continuous-time multiagent systems. In particular, we contribute with two novel solutions for this problem in different settings. First, we design distributed controllers that achieve synchronization in networked systems with homogeneous dynamics. The obtained controllers use time invariant feedback gains, and are static in the sense that no dynamic variables are required for control. Then, we solve the data-based synchronization problem for heterogeneous systems. This solution uses dynamic controllers that, different from the result in [12], do not require any knowledge of the systems’ models. To obtain these results, we provide a novel data-based representation of the synchronization error dynamics, leveraging the results in [6]. Different from the learning techniques in the literature [14, 15], we provide a method to collect persistently exciting data that guarantees the feasibility of the proposed solutions.
The paper is organized as follows. We first describe relevant results from the literature in Section II. Our solution to the synchronization problem in homogeneous multiagent systems is presented in Section III, and our method for the case of heterogeneous systems is shown in Section IV. We present a simulation example in Section V, and conclude the paper in Section VI.
@@ II-B Data-based representation and control of continuous-time systems
where $x\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{p}$ are the state, input and output vectors, respectively. Assume that input and output trajectories of (2) are measured, and the matrices $\mathcal{H}_{T}(u(t))$ and $\mathcal{H}_{T}(y(t))$ are defined as in (1). In some cases, we assume that state and state derivative information is also available, with the corresponding matrices $\mathcal{H}_{T}(x(t))$ and $\mathcal{H}_{T}(\dot{x}(t))$. The assumption that $\dot{x}$ can be either measured or estimated has been made in several works in the literature [17, 18]. This assumption can be relaxed by integrating the measured input-state trajectories instead (see, e.g., [19, Appendix A] and [20]). For simplicity of presentation, in this paper we let the trajectory $\dot{x}$ be available.
For data-based system representation and control, it is useful to collect persistently excited data from the CT system. A class of persistently exciting inputs for (2) is defined as follows.
@@
where $\alpha(t)\in\mathbb{R}^{M}$ satisfies
@@
$\left[\begin{array}[]{c}\mathcal{H}_{T}(u)\\ \mathcal{H}_{T}(x(0))\end{array}\right]\alpha(0)=\left[\begin{array}[]{c}\bar{u}(0)\\ \bar{x}(0)\end{array}\right],$ (7)
for $0\leq t<T$.
A straightforward consecuence of Theorem 1 is that the state trajectory $\bar{x}$ that corresponds to $(\bar{u},\bar{y})$ (the state realization being given by the data $x$ in (6)) can be written as
@@
We can also represent the derivative $\dot{\bar{x}}$ as follows. Taking the time derivative of (8), and noticing that $\mathcal{H}_{T}(x(t))\dot{\alpha}(t)=0$ as per (6), we obtain
@@
We conclude this subsection by commenting on control design. Different data-based controllers for CT systems have been proposed [9, 10]. A simple stabilizing controller was described in [21, Remark 2] as follows. Determine a matrix $\Lambda\in\mathbb{R}^{M\times n}$ such that
@@
$\mathcal{H}_{T}(\dot{x}(0))\Lambda+\Lambda^{\top}\mathcal{H}_{T}(\dot{x}(0))^{\top}\prec 0.$ (12)
Then, the matrix $K=-\mathcal{H}_{T}(u(0))\Lambda\left(\mathcal{H}_{T}(x(0))\Lambda\right)^{-1}$ is such that $A-BK$ is Hurwitz, i.e., system (2) is stabilized.
@@ II-C Graph theory definitions
Consider a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$ is a set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is a set of edges. In this paper, the nodes represent a set of $N$ dynamical agents and the edges correspond to their communication links. Associated with the graph edges, define the graph weights $a_{ij}$, $i,j=1,\ldots,N$, with $a_{ij}>0$ if $(v_{j},v_{i})\in\mathcal{E}$ and $a_{ij}=0$ otherwise. The graph is assumed to have no self-loops, i.e., $a_{ii}=0$ for all agents $i$. The adjacency matrix $\mathcal{A}\in\mathbb{R}^{N\times N}$ of the graph $\mathcal{G}$ is $\mathcal{A}=[a_{ij}]$. The weighted in-degree of each node $i$ is given by $d_{i}=\sum_{j=1}^{N}a_{ij}$, and the in-degree matrix of $\mathcal{G}$ is $D=\operatorname*{diag}_{i}\{d_{i}\}$. The Laplacian matrix of the graph is finally given by $\mathcal{L}=D-\mathcal{A}$.
Additional to the $N$ agents described above, consider an additional node, $v_{0}$, regarded as the leader node. Let $g_{i}$ be the pinning gain that represents the communication link between agent $i$ and the leader node. If agent $i$ can access the leader information, then $g_{i}>0$; otherwise $g_{i}=0$. The pinning matrix is defined as $G=\operatorname*{diag}_{i}\{g_{i}\}$. For convenience, we also defined the pinning vector $g\in\mathbb{R}^{N}$ as $g=[g_{1}\cdots g_{N}]^{\top}$.
@@ II-D The synchronization problem for homogeneous agents
where $x_{i}\in\mathbb{R}^{n}$ and $u_{i}\in\mathbb{R}^{m}$ are the state and input vectors of agent $i$, respectively. The pair $(A,B)$ is assumed to be controllable. The $N$ agents have limited sensing capabilities and can only access their own state information and that of their close neighbors as described by a communication graph $\mathcal{G}$. Moreover, let the leader dynamics be
@@
where $x_{0}\in\mathbb{R}^{n}$ is the leader state.
@@ II-E The synchronization problem for heterogeneous agents
Consider the case of a multiagent system with heterogeneous dynamics as
@@
where the state and input dimensions may differ between the systems, $x_{i}\in\mathbb{R}^{n_{i}}$, $u_{i}\in\mathbb{R}^{m_{i}}$. Every pair $(A_{i},B_{i})$ is assumed to be controllable. The output dimensions are assumed to be the same for all agents, $y_{i}\in\mathbb{R}^{p}$, $i=1,\ldots,N$. Similarly, the leader has the following dynamics
@@
Although the systems (15)-(16) have outputs $y_{i}$, it is a common assumption in the synchronization literature that the complete state $x_{i}$ is measurable (see, e.g., [2, 12]). Rather than measurements, the outputs $y_{i}$ represent the signals to be synchronized. If only output measurements are available, as in, e.g., [3], an observer can be included in the control loop. We do not consider this setting in this paper.
@@
We can now present our data-based solutions for Problems 1 and 2. Our reason for offering different solutions to these problems is that, while we solve Problem 2 using dynamic controllers, our solution to Problem 1 does not require the use of dynamic variables for control. This latter case is described in the following section.
@@ III-A A result on model-based synchronization control
where
@@
$u=[u_{1}^{\top}\quad u_{2}^{\top}\;\cdots\;u_{N}^{\top}]^{\top}\in\mathbb{R}^{mN},$ (21)
and $\otimes$ represents the Kronecker product.
@@
$K=c\bar{K},$ (24)
where $c$ is a large enough scalar, and
$\bar{K}=R^{-1}B^{\top}\bar{P}$ (25)
with $\bar{P}\in\mathbb{R}^{n\times n}$ being the solution of the algebraic Riccati equation (ARE)
@@
and where $Q\in\mathbb{R}^{n\times n}$ and $R\in\mathbb{R}^{m\times m}$ are user-defined positive definite matrices.
@@
Consider the multiagent system (13)-(14), let the pair $(A,B)$ be controllable, and let Assumption 1 hold. There exists a matrix $K\in\mathbb{R}^{m\times n}$ as in (24) and a block diagonal matrix $P=\operatorname*{blkdiag}_{i}\{P_{i}\}$, $P_{i}\in\mathbb{R}^{n\times n}$, $i=1,\ldots,N$, such that $P\succ 0$ and (27) holds with $A_{c}$ as in (23).
@@
It was shown in [22, Theorem 1] that, under Assumption 1, there exists a diagonal matrix $S=\operatorname*{diag}_{i}\{s_{i}\}$, $s_{i}>0$, $i=1,\ldots,N$, such that $(\mathcal{L}+G)S+S(\mathcal{L}+G)=\bar{Q}\succ 0$. Observe that, since $\bar{Q}\succ 0$, there exists a small enough constant $\epsilon>0$ such that $\bar{Q}-\epsilon I\succ 0$.
@@ III-B Data-based representation of the synchronization error dynamics
for $0\leq t<T$. Notice that any leader trajectory $\bar{x}_{0}$ can also be obtained as
$\mathcal{H}_{T}(x(t))\alpha_{0}(t)=\bar{x}_{0}(t),$
where $\alpha_{0}$ also satisfies
@@
$\delta_{i}(t)=$ $\sum_{j=1}^{N}a_{ij}\mathcal{H}_{T}(x(t))\left(\alpha_{i}(t)-\alpha_{j}(t)\right)$
$\qquad\qquad+g_{i}\mathcal{H}_{T}(x(t))\left(\alpha_{i}(t)-\alpha_{0}(t)\right)$
$=$ $\mathcal{H}_{T}(x(t))\Bigl{[}\sum_{j=1}^{N}a_{ij}\left(\alpha_{i}(t)-\alpha_{j}(t)\right)$
$\qquad\qquad+g_{i}\left(\alpha_{i}(t)-\alpha_{0}(t)\right)\Bigr{]}.$ (30)
@@
For the dynamics of the synchronization error $\delta$, from (31) and the fact that $\mathcal{H}_{T}(x(t))\dot{\alpha}_{i}(t)=0$ for all $i=0,\ldots,N$ from (6), we obtain
@@ III-C Stabilization of the synchronization error
Taking advantage of the representations in (31) and (32), we can use known methods for data-based control to stabilize the synchronization error. In the following, we make use of the procedure in (11)-(12), although any other of the methods in the literature can potentially be employed.
We begin by showing the following lemma. For convenience, define the matrices $\mathcal{L}_{g}=[-g\quad\mathcal{L}+G]$ and $\mathcal{I}_{g}=[0_{N\times 1}\quad I_{N}]$.
@@
A direct application of (11)-(12) to the data-based representations (31) and (32) provides a stabilizing global controller, as we show in the following lemma. However, this does not correspond to a proper solution of Problem 1 due to the fact that, in general, distributed controllers as in (22) are not obtained.
@@
The problem (35) is feasible. Moreover, the input $u=-K\delta$ with $K=-(\mathcal{I}_{g}\otimes\mathcal{H}_{T}(u))\Lambda((\mathcal{L}_{g}\otimes\mathcal{H}_{T}(x(0)))\Lambda)^{-1}$ renders the global synchronization error dynamics (19) asymptotically stable.
@@
Notice that each $\delta^{k}\in\mathbb{R}^{nN}$ can be expressed in data-based fashion as in (31), i.e., $\delta^{k}=(\mathcal{L}_{g}\otimes\mathcal{H}_{T}(x(0)))\alpha^{k}$ with $\alpha^{k}\in\mathbb{R}^{M(N+1)}$, and similarly for $\dot{\delta}^{k}$ using (32). Defining the matrix $\Lambda:=\left[\alpha^{1}\quad\alpha^{2}\;\cdots\;\alpha^{nN}\right]\in\mathbb{R}^{M(N+1)\times nN}$, we can write
@@
Finally, we show that it is always possible to select $\Lambda$ such that
@@
Lemma 4 provides a global input of the form $u=-K\delta$ which does not necessarily correspond to the distributed inputs (22) required in Problem 1. In the following theorem, we show that this issue can be solved with additional constraints in the LMIs (11)-(12).
@@
The problem (38) is feasible. Moreover, the distributed inputs (22) with matrices $K_{i}=F_{i}P_{i}^{-1}$, $i=1,\ldots,N$, solve Problem 1.
@@
From Lemma 4 it follows that the conditions (38a)-(38d) provide a stabilizing global input of the form $u=-K\delta$, where $K=-(\mathcal{I}_{g}\otimes\mathcal{H}_{T}(u))\Lambda((\mathcal{L}_{g}\otimes\mathcal{H}_{T}(x))\Lambda)^{-1}$. The conditions (38a) and (38e) constrain the matrices to be block diagonal, such that $K=-\operatorname*{blkdiag}_{i}\{F_{i}\}\operatorname*{blkdiag}_{i}\{P_{i}^{-1}\}=\operatorname*{blkdiag}_{i}\{K_{i}\}$ with $K_{i}$ as in the theorem statement.
@@
Theorem 2 provides a data-based solution to Problem 1, which considers groups of identical agents. In the following section, we address the case of heterogeneous agents as described in Problem 2.

### v2

@@ Abstract
This paper presents novel solutions of the data-based synchronization problem for continuous-time multiagent systems. We consider the cases of homogeneous and heterogeneous systems. First, we obtain a data-based representation of the synchronization error dynamics for homogeneous systems and show how to extend existing data-based stabilization results to stabilize such error dynamics. The proposed method relies on the solution of a set of linear matrix inequalities that are shown to be feasible. Then, we solve the synchronization problem for heterogeneous systems by means of dynamic controllers. Different from existing results, we do not require model knowledge for the followers and the leader. The theoretical results are finally validated using a numerical simulation.
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I INTRODUCTION
In recent years, there has been a great interest in developing stabilizing controllers without the need of any knowledge of a system’s model. One of the main drivers of this recent research has been Willems’ fundamental lemma [5], which provides a method to represent every input-output trajectory of a discrete-time (DT) linear system using only measured data. Continuous-time (CT) versions of this result have also been proposed [6, 7]. Willems’ lemma has been recently leveraged for the data-based control of DT (see the survey [8]) and CT systems (see, e.g., [9, 10, 11]).
Some results have been obtained for the data-driven control of multiagent systems. In [12, 13] the synchronization problem is solved using only partial knowledge of the agents’ dynamics. In [14], consensus protocols are designed for systems with identical models using time-varying feedback matrices. Several iterative learning algorithms have been proposed to achieve synchronization as in, e.g., [15, 16]. In [17], a multiagent predictive control scheme is used to minimize a global cost function.
In this paper, we study the data-based synchronization problem for continuous-time multiagent systems. In particular, we contribute with two novel solutions for this problem in different settings. First, we design distributed controllers that achieve synchronization in networked systems with homogeneous dynamics. The obtained controllers use time invariant feedback gains, and are static in the sense that no dynamic variables are required for control. Then, we solve the data-based synchronization problem for heterogeneous systems. This solution uses dynamic controllers that, different from the result in [12, 13], do not require any knowledge of the systems’ models. To obtain these results, we provide a novel data-based representation of the synchronization error dynamics, leveraging the results in [6]. Different from the learning techniques in the literature [15, 16], we provide a method to collect persistently exciting data that guarantees the feasibility of the proposed solutions.
The paper is organized as follows. Relevant results from the literature are described in Section II. Our solution to the synchronization problem in homogeneous systems is presented in Section III, and our method for the heterogeneous case is shown in Section IV. We present a simulation example in Section V, and conclude the paper in Section VI.
@@ II-B Data-based representation and control of continuous-time systems
where $x\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{p}$ are the state, input and output vectors, respectively. Assume that input and output trajectories of (2) are measured, and the matrices $\mathcal{H}_{T}(u(t))$ and $\mathcal{H}_{T}(y(t))$ are defined as in (1). In some cases, we assume that state and state derivative information is also available, with the corresponding matrices $\mathcal{H}_{T}(x(t))$ and $\mathcal{H}_{T}(\dot{x}(t))$. The assumption that $\dot{x}$ can be either measured or estimated has been made in several works in the literature [18, 19]. This assumption can be relaxed by integrating the measured input-state trajectories instead (see, e.g., [20, Appendix A] and [21]). For simplicity of presentation, in this paper we let the trajectory $\dot{x}$ be available.
For data-based system representation and control, it is useful to collect persistently excited (PE) data from the CT system. A class of PE inputs for (2) is defined as follows.
@@
for $0\leq t<T$, where $\alpha(t)\in\mathbb{R}^{M}$ satisfies
@@
$\left[\begin{array}[]{c}\mathcal{H}_{T}(u)\\ \mathcal{H}_{T}(x(0))\end{array}\right]\alpha(0)=\left[\begin{array}[]{c}\bar{u}(0)\\ \bar{x}(0)\end{array}\right].$ (7)
A straightforward consequence of Theorem 1 is that the state trajectory $\bar{x}$ that corresponds to $(\bar{u},\bar{y})$ (the state realization is given by the data $x$ in (6)) can be written as
@@
Taking the time derivative of (8), and noticing that $\mathcal{H}_{T}(x(t))\dot{\alpha}(t)=0$ as per (6), we also obtain
@@
We conclude this subsection by commenting on control design. Different data-based controllers for CT systems have been proposed [9, 10]. A simple stabilizing controller was described in [22, Remark 2] as follows. The matrix $K=-\mathcal{H}_{T}(u(0))\Lambda\left(\mathcal{H}_{T}(x(0))\Lambda\right)^{-1}$, where $\Lambda\in\mathbb{R}^{M\times n}$ satisfies
@@
$\mathcal{H}_{T}(\dot{x}(0))\Lambda+\Lambda^{\top}\mathcal{H}_{T}(\dot{x}(0))^{\top}\prec 0,$ (12)
is such that $A-BK$ is Hurwitz, i.e., system (2) is stabilized.
@@ II-C Graph theory definitions
Consider a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$ is a set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is a set of edges. In this paper, the nodes represent a set of $N$ dynamical agents and the edges correspond to their communication links. Define the graph weights $a_{ij}$, $i,j=1,\ldots,N$, with $a_{ij}>0$ if $(v_{j},v_{i})\in\mathcal{E}$ and $a_{ij}=0$ otherwise. The graph is assumed to have no self-loops, i.e., $a_{ii}=0$ for all agents $i$. The adjacency matrix $\mathcal{A}\in\mathbb{R}^{N\times N}$ of the graph $\mathcal{G}$ is $\mathcal{A}=[a_{ij}]$. The weighted in-degree of each node $i$ is given by $d_{i}=\sum_{j=1}^{N}a_{ij}$, and the in-degree matrix of $\mathcal{G}$ is $D=\operatorname*{diag}_{i}\{d_{i}\}$. The Laplacian matrix of the graph is finally given by $\mathcal{L}=D-\mathcal{A}$.
Additional to the $N$ agents described above, consider an additional node, $v_{0}$, regarded as the leader node. The pinning gain $g_{i}$ is such that $g_{i}>0$ if agent $i$ can access the leader information; otherwise $g_{i}=0$. The pinning matrix is defined as $G=\operatorname*{diag}_{i}\{g_{i}\}$. For convenience, we also defined the pinning vector $g\in\mathbb{R}^{N}$ as $g=[g_{1}\cdots g_{N}]^{\top}$.
@@ II-D The synchronization problem for homogeneous agents
where $x_{i}\in\mathbb{R}^{n}$ and $u_{i}\in\mathbb{R}^{m}$ are the state and input vectors of agent $i$, respectively. The leader dynamics are given by
@@
where $x_{0}\in\mathbb{R}^{n}$ is the leader state. The $N$ agents can only access their own state information and that of their close neighbors as described by a communication graph $\mathcal{G}$.
@@ II-E The synchronization problem for heterogeneous agents
Consider a heterogeneous multiagent system as
@@
where the state and input dimensions may differ between the systems, $x_{i}\in\mathbb{R}^{n_{i}}$, $u_{i}\in\mathbb{R}^{m_{i}}$. The output dimensions are the same for all agents, $y_{i}\in\mathbb{R}^{p}$. The leader has the dynamics
@@
Although the systems (15)-(16) have outputs $y_{i}$, it is a common assumption in the synchronization literature that the complete state $x_{i}$ is measurable (see, e.g., [2, 12]). Rather than measurements, the outputs $y_{i}$ represent the signals to be synchronized. If only output measurements are available as in, e.g., [3], an observer can be included in the control loop. We do not consider this setting in this paper.
@@
We now present our data-based solutions for Problems 1 and 2. Our reason for solving these problems separately is that, different from our solution to Problem 2, we solve Problem 1 without the use of dynamic variables for control. This latter case is described in the following section.
@@ III-A A result on model-based synchronization control
where $\otimes$ represents the Kronecker product, and
@@
$u=[u_{1}^{\top}\quad u_{2}^{\top}\;\cdots\;u_{N}^{\top}]^{\top}\in\mathbb{R}^{mN}.$ (21)
@@
$K=cR^{-1}B^{\top}\bar{P},$ (24)
where $c$ is a large enough scalar, $\bar{P}\in\mathbb{R}^{n\times n}$ is the solution of the algebraic Riccati equation (ARE)
@@
and $Q\in\mathbb{R}^{n\times n}$ and $R\in\mathbb{R}^{m\times m}$ are user-defined positive definite matrices.
@@
Consider the multiagent system (13)-(14), let the pair $(A,B)$ be controllable, and let Assumption 1 hold. There exists a matrix $K\in\mathbb{R}^{m\times n}$ and a block diagonal matrix $P=\operatorname*{blkdiag}_{i}\{P_{i}\}$, $P_{i}\in\mathbb{R}^{n\times n}$, $i=1,\ldots,N$, such that $P\succ 0$ and (26) holds with $A_{c}$ as in (23).
@@
It was shown in [23, Theorem 1] that, under Assumption 1, there exists a diagonal matrix $S=\operatorname*{diag}_{i}\{s_{i}\}$, $s_{i}>0$, $i=1,\ldots,N$, such that $(\mathcal{L}+G)S+S(\mathcal{L}+G)=\bar{Q}\succ 0$. Observe that, since $\bar{Q}\succ 0$, there exists a small enough constant $\epsilon>0$ such that $\bar{Q}-\epsilon I\succ 0$.
@@ III-B Data-based representation of the synchronization error dynamics
for $0\leq t<T$. Any leader trajectory $\bar{x}_{0}$ can also be obtained as $\mathcal{H}_{T}(x(t))\alpha_{0}(t)=\bar{x}_{0}(t)$, where $\alpha_{0}$ also satisfies
@@
$\delta_{i}(t)=\mathcal{H}_{T}(x(t))\Bigl{[}\sum_{j=1}^{N}a_{ij}\left(\alpha_{i}(t)-\alpha_{j}(t)\right)\\ +g_{i}\left(\alpha_{i}(t)-\alpha_{0}(t)\right)\Bigr{]}.$ (29)
@@
For the dynamics of $\delta$, from (30) and the fact that $\mathcal{H}_{T}(x(t))\dot{\alpha}_{i}(t)=0$ for all $i=0,\ldots,N$ from (6), we obtain
@@ III-C Stabilization of the synchronization error
Taking advantage of the representations in (30) and (31), we can use any data-based controller in the literature to stabilize the synchronization error. In the following, we use the procedure in (11)-(12). We begin by showing the following lemma. For convenience, define the matrices $\mathcal{L}_{g}=[-g\quad\mathcal{L}+G]$ and $\mathcal{I}_{g}=[0_{N\times 1}\quad I_{N}]$.
@@
A direct application of (11)-(12) to the data-based representations (30) and (31) provides a stabilizing global controller, as we show in the following lemma. However, this is not a proper solution to Problem 1 because, in general, distributed controllers as in (22) are not obtained.
@@
The problem (34) is feasible. Moreover, the input $u=-K\delta$ with $K=-(\mathcal{I}_{g}\otimes\mathcal{H}_{T}(u))\Lambda((\mathcal{L}_{g}\otimes\mathcal{H}_{T}(x(0)))\Lambda)^{-1}$ renders the global error dynamics (19) asymptotically stable.
@@
Notice that each $\delta^{k}\in\mathbb{R}^{nN}$ can be expressed in data-based fashion as in (30), i.e., $\delta^{k}=(\mathcal{L}_{g}\otimes\mathcal{H}_{T}(x(0)))\alpha^{k}$ with $\alpha^{k}\in\mathbb{R}^{M(N+1)}$, and similarly for $\dot{\delta}^{k}$ using (31). Defining $\Lambda:=\left[\alpha^{1}\quad\alpha^{2}\;\cdots\;\alpha^{nN}\right]\in\mathbb{R}^{M(N+1)\times nN}$, we can write
@@
Finally, we show that we can always select $\Lambda$ such that
@@
Lemma 4 provides a global input of the form $u=-K\delta$ which does not necessarily correspond to the distributed inputs required in Problem 1 (see, e.g., (22)). In the following theorem, we show that this issue can be solved with additional constraints in the LMIs (34).
@@
The problem (37) is feasible. Moreover, the distributed inputs $u_{i}=-K_{i}\delta_{i}$ with $K_{i}=F_{i}P_{i}^{-1}$, $i=1,\ldots,N$, solve Problem 1.
@@
From Lemma 4 it follows that (37a)-(37d) provide a stabilizing global input of the form $u=-K\delta$, where $K=-(\mathcal{I}_{g}\otimes\mathcal{H}_{T}(u))\Lambda((\mathcal{L}_{g}\otimes\mathcal{H}_{T}(x(0)))\Lambda)^{-1}$. The conditions (37a) and (37e) constrain the matrices to be block diagonal, such that $K=-\operatorname*{blkdiag}_{i}\{F_{i}\}\operatorname*{blkdiag}_{i}\{P_{i}^{-1}\}=\operatorname*{blkdiag}_{i}\{K_{i}\}$ with $K_{i}$ as in the theorem statement.
@@
In the following section, we address the synchronization problem for heterogeneous agents as described in Problem 2.